Calculate or hydrolysis a 0.1 KCN solution at 25 oC. The HCN ionization constant is 7.0 x 10-10 and the ionic product of water is 1.0 x 10-14.
1
Expert's answer
2020-12-16T05:54:26-0500

"CN^- + H_2O \quad^\to _\leftarrow \quad HCN + OH^-"

before reaction

"C + - \quad^\to _\leftarrow \quad - + \ -"

after reaction

"C(1-h) + - \quad^\to _\leftarrow \quad Ch + Ch"


"K_h = \dfrac{[Ch][Ch]}{[C(1-h)]}"


Let 'h << 1

So, "K_h = \dfrac{C^2 h ^2}{C}"


"K_h = Ch^2"


We know that for a salt of weak acid and strong base;


"K_h =\dfrac{K_w}{K_a} =\dfrac{1×10^{-14}}{7×10^{-10}} = 1.43×10^{-5}"


"Ch^2 = 1.43\u00d710^{-5}"


"0.1h^2 = 1.43\u00d710^{-5}"


"h^2 = 1.43\u00d710^{-4}"


h = 0.012